I am having some trouble solving this problem because I forgot how to set it up. I just need the first two steps and I can figure it out from there. I need to solve it using a logarithmic equation: 3^(2x+7)=3^(1/2)
Answer by stanbon(57347)

3^(2x+7)=3^(1/2)
---
Take the log(base 3) of both sides to get:
2x+7 = 1/2
Then solve for "x".
=======================
Note: log(base a) of a^x = x
=======================
Cheers,
Stan H.